OpenStudy (anonymous):

LAST QUESTION! Write the equation in general format for the line passing through (2,-6) and (8,3)

OpenStudy (anonymous):

y=mx+b u could find m ( slope ) and b

OpenStudy (lgbasallote):

im just going to write the two-point formula don't mind me.... \(\LARGE y - y{1} = (\frac{y_{2} - y_{1}}{x_{2} - x_{1}}) (x - x_{1})\)

OpenStudy (lgbasallote):

that's supposed to be \(y - y_{1}\)

OpenStudy (anonymous):